# Is the System Message Really Important to Jailbreaks in Large Language Models?

### v1

@@ Abstract
The rapid evolution of Large Language Models (LLMs) has rendered them indispensable in modern society. While security measures are typically in place to align LLMs with human values prior to release, recent studies have unveiled a concerning phenomenon named ”jailbreak.” This term refers to the unexpected and potentially harmful responses generated by LLMs when prompted with malicious questions. Existing research focuses on generating jailbreak prompts but our study aim to answer a different question: Is the system message really important to jailbreak in LLMs? To address this question, we conducted experiments in a stable GPT version gpt-3.5-turbo-0613 to generated jailbreak prompts with varying system messages: short, long, and none. We discover that different system messages have distinct resistances to jailbreak by experiments. Additionally, we explore the transferability of jailbreak across LLMs. This finding underscores the significant impact system messages can have on mitigating LLMs jailbreak. To generate system messages that are more resistant to jailbreak prompts, we propose System Messages Evolutionary Algorithms (SMEA). Through SMEA, we can get robust system messages population that demonstrate up to 98.9% resistance against jailbreak prompts. Our research not only bolsters LLMs security but also raises the bar for jailbreak, fostering advancements in this field of study.
Warning: Some of the examples may be harmful!
Machine Learning, ICML
@@ 1 Introduction
The rapid evolution of Large Language Models (LLMs), including prominent examples such as ChatGPT (OpenAI, 2022), Bard (Google, 2023a), and Claude (Google, 2023a), has ushered in an era where conversing with intelligent systems possessing human-like language abilities is feasible. However, the training data utilized for these LLMs is typically from the internet, inevitably incorporating malicious content. This poses a significant risk, potentially leading LLMs to produce offensive output (Hazell, 2023; Kang et al., 2023). In response, various alignment methods, such as Reinforcement Learning through Human Feedback (RLHF) (Bai et al., 2022a; Ouyang et al., 2022; Christiano et al., 2017; Bai et al., 2022b), are imposed prior to the release LLMs. These measures aim to ensure that the outputs generated by LLMs are both useful and harmless, mitigating the potential harm caused by offensive content.
@@
However, the LLMs community like JailbreakChat (Albert, 2023), continues to witness the emergence of various queries capable of prompting LLMs to generate malicious content. This behavior is termed ”jailbreak” of LLMs (Wei et al., 2023). Jailbreak prompts, which trigger such behavior, have garnered significant attention in recent research endeavors (Perez et al., 2022; Ganguli et al., 2022; Shen et al., 2023). As depicted in Figure 1, when a direct query just involves only a harmful question, the LLMs will reject answering it. However, when we query a jailbreak prompt including a harmful question, the LLMs will respond affirmatively to the question. Initially, researchers manually crafted queries to gradually steer LLMs towards jailbreak, exemplified by the famous ”grandmother loophole” tactic (Reddit, 2023), exploiting familial relationships to manipulate LLMs behavior. However, manual crafting of jailbreak prompts is inefficient, prompting a shift towards developing systematic and automated methods for generating such prompts (Zou et al., 2023; Huang et al., 2023; Lapid et al., 2023; Carlini et al., 2023; Qi et al., 2023; Chao et al., 2023; Yu et al., 2023; Li et al., 2023).
Figure 1: An example of jailbreak prompt in gpt-3.5-turbo-0613
Despite the focus on generating jailbreak prompts, an overlooked aspect is the role of system messages. System messages (OpenAI, 2022), integral for developers interfacing with LLMs, have not received systematic investigation concerning their influence on jailbreak. Typically, system messages are set like ”You are a helpful, respectful, and honest assistant.” (Huggingface, 2023b) but we are not sure what the system messages do. Consequently, our study aims to address the follwing questions:
Is the system message indeed crucial to jailbreaks in LLMs? And if so, how can we optimize system messages to bolster resistance against such jailbreak prompts?
Through systematic exploration, we seek to shed light on this overlooked aspect of LLMs security, paving the way for more effective mitigation strategies against jailbreak.
@@
To answer the first question, it is imperative to conduct extensive test on LLMs with various system messages. Our comprehensive experimentation unequivocally demonstrates the crucial role of system messages in jailbreaks of LLMs. Just setting the system message to ”You are a helpful, respectful, and honest assistant.” will make jailbreak easy. For promoting the security of LLMs, we design the System Messages Evolutionary Algorithm(SMEA) to search the diverse and more resistant system messages to jailbreak. This endeavor holds immense promise in enhancing LLMs security, empowering service providers to embed carefully crafted, inaccessible system messages within LLMs as a preventive measure against jailbreak exploits.
To sum up, the main contributions of this paper are as follows:
@@
  • We introduce a novel perspective by examining the role of system messages in LLMs security and jailbreak research.
  • We first conduct comprehensive tests on the impact of system messages using gpt-3.5-turbo-0613, exploring the question: ”Is the system message really important to jailbreaks in LLMs?” We discover that varying system messages can greatly influence the transferability of jailbreak prompts. This phenomenon affirms the critical significance of system messages in facilitating or mitigating jailbreak. Furthermore, we demonstrate that LLMs equipped with robust system messages exhibit enhanced security.
@@
  • To bolster resistance against LLMs jailbreak prompts, we delve into the optimization of system messages. We use evolution algorithms because the truth that most LLMs are black box models. Specifically, we propose two System Message Evolution Algorithms (SMEA): SMEA-R and SMEA-C, each tailored to generate diverse and highly secure system messages. Through testing, we validate the efficacy of SMEA, demonstrating its robustness in resisting up to 98.9% of jailbreak prompts. This not only fortifies LLMs security but also elevates the threshold for successful jailbreak prompts, thereby advancing the research landscape in LLMs security.
The structure of the remainder of the paper is outlined as follows. We commence with an exposition of background knowledge in Section 2, providing essential context for the subsequent discussions. Following this, in Section 3, we present the datasets, model specifications, and evaluation criteria utilized in our study. Moving forward to Section 4, we delve into the generation of jailbreak prompts tailored for gpt-3.5-turbo-0613, exploring the impact of different system messages on model vulnerability. Additionally, we investigate the transferability of jailbreak prompts across varied system message configurations within gpt-3.5-turbo-0613. In Section 5, we introduce the SMEA algorithm, designed to optimize system prompts for enhanced resilience against jailbreak prompts. Subsequently, in Section 6, we present experimental results showcasing the efficacy of the SMEA algorithm in bolstering LLMs security. Finally, we encapsulate our findings and contributions in Section 7, providing a comprehensive summary of the paper’s key insights and implications.
@@ 2 Background Knowledge
In this section, we aim to offer a comprehensive overview of the background knowledge and pertinent terminology to this paper’s discussion. We begin by introducing related information about LLMs. Furthermore, we delve into the concept of system messages within the realm of LLMs. Subsequently, we provide a detailed exposition on jailbreak prompts, shedding light on their significance and implications. Lastly, given the incorporation of evolutionary algorithms in our study, we provide a general overview of how evolutionary algorithms leverage populations to iteratively obtain high-quality solutions.
@@ 2.1 Large Language Models
Large Language Models (LLMs), characterized by their extensive parameter count and training on vast corpora of human natural language, have emerged as a cornerstone in deep learning frameworks (Chang et al., 2023). With parameter counts reaching into the billions, LLMs possess language capabilities akin to human fluency.
Presently, mainstream LLMs such as ChatGPT (OpenAI, 2022), GPT4 (Bubeck et al., 2023), Llama2 (Touvron et al., 2023) are founded upon the transformer decoder architecture, operating in an autoregressive manner. This operational mode enables these models to predict subsequent words in a sequence based on preceding input text. Upon receiving input $\{w_{1},w_{2},\cdots,w_{t}\}$, LLMs forecast the most probable next word, $w_{t+1}$, subsequently incorporating it into the input sequence $\{w_{1},w_{2},\cdots,w_{t},w_{t+1}\}$ for further prediction. This iterative process continues until an end-of-sequence token is reached or the maximum sequence length supported by the LLMs is attained.
In developer mode, when interfacing with LLMs via the API, accessing to system messages is facilitated. These system messages typically exist in both long and short formats, as outlined in Table 4. Conventionally, system messages serve as a protective barrier to reinforce system security. However, despite their presumed significance, systematic investigations into system messages remain scarce. Hence, our research endeavors to bridge this gap by conducting a comprehensive exploration of system messages.
@@ 2.2 Jailbreak prompts
Jailbreak prompts are meticulously constructed inputs engineered to provoke LLMs into generating harmful content (Wei et al., 2023). Illustrated in Figure 1, we provide an example of a jailbreak prompt capable of coaxing the gpt-3.5-turbo-0613 model to produce deleterious output, with the red segments representing malicious queries. Despite concerted efforts to bolster LLMs’ resistance against such prompts (Bai et al., 2022a; Ouyang et al., 2022; Christiano et al., 2017; Bai et al., 2022b), existing literature suggests persistent vulnerabilities (Zou et al., 2023; Huang et al., 2023; Lapid et al., 2023; Carlini et al., 2023; Qi et al., 2023; Chao et al., 2023; Yu et al., 2023; Li et al., 2023). For open source LLM, Zou et al. (2023) proposed a novel jailbreaking approach by appending suffixes to malicious questions, leveraging the autoprompt (Shin et al., 2020) concept for open-source models, while also elucidating the transferability of jailbreak prompts. Huang et al. (2023) underscored the profound impact of parameter configurations on LLMs vulnerabilities. For black box LLMs, Lapid et al. (2023) employed genetic algorithms to craft adversarial suffixes for breaching black-box LLMs. Moreover, Yu et al. (2023) advocated for applying software testing principles to LLMs jailbreaking, while Chao et al. (2023) proposed using one LLM to infiltrate another. Li et al. (2023) engineered a multi-step jailbreak attack against ChatGPT, targeting the extraction of sensitive personal data, thereby raising profound privacy concerns and significant security risks for LLMs applications. In addition to the above recent investigations, Carlini et al. (2023) have demonstrated that simply manipulating image inputs can compromise the security of multi-modal models. Furthermore, research (Qi et al., 2023) suggests that fine-tuning aligned LLMs on adversarial examples can jeopardize their secure alignment. There is no doubt that these prompts pose a significant obstacle to the widespread use of LLMs.
@@ 2.3 Evolutionary Algorithms
Evolutionary algorithms (Eiben et al., 2015), renowned as heuristic search techniques, using the principle of natural selection to optimize populations. Prominent algorithms in this domain include NSGA-II (Deb et al., 2002), IBEA (Zitzler & Künzli, 2004), and MOEA/D (Zhang & Li, 2007), all of which encompass three fundamental operators: namely crossover, mutation, and selection.
@@
    Crossover operator: This operator selects two individuals from the population and combines their information to produce offsprings. This process is repeated until the desired number of offspring is generated, maintaining the population size.
@@
    Selection operator: Following mutation, the new population is merged with the previous population. Individuals are then selected based on their fitness, with the best-performing individuals being chosen to form the new population for the next iteration.
Through successive iterations of these operations, individuals with inferior performance are gradually replaced, leading to the emergence of optimal and diverse individuals within the final population.
@@ 3.1 Models and Datasets
We select gpt-3.5-turbo-0613 (OpenAI, 2022) as our stable version to investigate the impact of system messages on jailbreaks of LLMs. It’s worth noting that the gpt-3.5-turbo-0613 chat model has undergone iterative enhancements through the application of RLHF (Bai et al., 2022a; Ouyang et al., 2022; Christiano et al., 2017; Bai et al., 2022b) techniques. To systematically explore this impact, we utilize GPTFuzzer (Yu et al., 2023) to generate jailbreak prompts from a collection of $77$ prompt templates and $100$ questions. Specifically, we employ GPTFuzzer to generate $300$ jailbreak prompts on gpt-3.5-turbo-0613 under three conditions respectively: with a short system message, a long system message, and no system message. Details regarding these system messages are provided in Appendix A Table 4.
@@ 3.2 Evaluation
In our study, we evaluate the success of LLMs jailbreak using the Attack Success Rate (ASR) metric. Previous studies have employed various criteria to evaluate the success of jailbreak, which we summarize as follows:

### v2

# Is the System Message Really Important for Jailbreaks in Large Language Models?

@@ Abstract
The rapid evolution of Large Language Models (LLMs) has rendered them indispensable in modern society. While security measures are typically to align LLMs with human values prior to release, recent studies have unveiled a concerning phenomenon named "Jailbreak". This term refers to the unexpected and potentially harmful responses generated by LLMs when prompted with malicious questions. Most existing research focus on generating jailbreak prompts but system message configurations vary significantly in experiments. In this paper, we aim to answer a question: Is the system message really important for jailbreaks in LLMs? We conduct experiments in mainstream LLMs to generate jailbreak prompts with varying system messages: short, long, and none. We discover that different system messages have distinct resistances to jailbreaks. Therefore, we explore the transferability of jailbreaks across LLMs with different system messages. Furthermore, we propose the System Messages Evolutionary Algorithm (SMEA) to generate system messages that are more resistant to jailbreak prompts, even with minor changes. Through SMEA, we get a robust system messages population with little change in the length of system messages. Our research not only bolsters LLMs security but also raises the bar for jailbreaks, fostering advancements in this field of study.
Is the System Message Really Important for Jailbreaks in Large Language Models?
Anonymous ACL submission
@@ 1 Introduction
The rapid evolution of Large Language Models (LLMs), including ChatGPT OpenAI (2022), Bard Google (2023a), and Claude Google (2023a), has ushered in an era where conversing with intelligent systems possessing human-like language abilities is feasible. However, the training data for these LLMs is typically from the internet, inevitably incorporating malicious content. This poses a significant risk, potentially leading LLMs to produce offensive output Hazell (2023); Kang et al. (2023). In response to the situation, various alignment methods, such as different Reinforcement Learning from Human Feedback (RLHFBai et al. (2022a); Ouyang et al. (2022); Christiano et al. (2017); Bai et al. (2022b), are imposed prior to the release of LLMs. These measures aim to ensure that the outputs generated by LLMs are both useful and harmless, reducing the likelihood of aggressive content appearing.
@@
Despite the efforts of researchers, the LLM community, such as JailbreakChat Albert (2023), continues to witness the emergence of various queries capable of prompting LLMs to generate malicious content. These behaviors are termed "jailbreak" of LLMs Wei et al. (2023); Liu et al. (2023). Jailbreak prompts, which trigger jailbreaks, have garnered significant attention in recent research Perez et al. (2022); Ganguli et al. (2022); Shen et al. (2023). As depicted in Figure LABEL:jail_example (a) (b), when a direct query just involves only a harmful question, GPT3.5-turbo-0613 OpenAI (2022) will reject answering it. However, when we query a jailbreak prompt including a harmful question, GPT3.5-turbo-0613 will respond affirmatively to the question. Initially, researchers manually crafted queries to gradually steer LLMs towards jailbreaks, exemplified by the famous "grandmother loophole" tactic Reddit (2023), exploiting familial relationships to manipulate LLMs behavior. However, manual crafting of jailbreak prompts is inefficient, prompting a shift towards developing systematic and automated methods for generating such prompts Zou et al. (2023); Huang et al. (2023); Lapid et al. (2023); Carlini et al. (2023); Qi et al. (2023); Chao et al. (2023); Yu et al. (2023); Li et al. (2023).
Most researchers have focused primarily on jailbreaking techniques, overlooking the potential impact of system messages on the effectiveness of these techniques. Hence, in previous worksZou et al. (2023); Huang et al. (2023); Lapid et al. (2023); Carlini et al. (2023); Qi et al. (2023); Chao et al. (2023); Yu et al. (2023); Li et al. (2023), researchers employed various configurations for system messages. Despite their importance for developers working with LLMs, system messages have not been systematically investigated in terms of their influence on jailbreaking attempts. System messages are often set to phrases like "You are a helpful, respectful, and honest assistant" Huggingface (2023b); however, it remains uncertain whether such messages are optimal for ensuring the security of LLMs in black box. Consequently, this study aims to address the following research questions:
RQ1: Is the system message really important for jailbreaks in LLMs? RQ2: If so, Whether the longer system system message is better at protecting LLMs? RQ3: Does jailbreak rate of LLMs change significantly when there is a little change (similar in semantics, a little change in length) in system messages? RQ4: How can we optimize system messages to enhance resistance against such jailbreak prompts?
@@
To address the first research question, it is essential to conduct extensive tests on LLMs using a diverse set of system messages. Our comprehensive experimentation unequivocally demonstrates the crucial role of system messages in jailbreaks of LLMs. Just setting the system message to "You are a helpful, respectful, and honest assistant." will make jailbreaks easy. By comparing the jailbreak rates of different system messages, we answer the second question. Furthermore, our experiments reveal that different LLMs exhibit significant variations in their responses to minor changes in system messages. Finally, for promoting the security of LLMs, we design the System Messages Evolutionary Algorithm (SMEA) to search for diverse and more resistant system messages to jailbreaks. This approach has significant potential to enhance LLM security by enabling service providers to embed meticulously designed, concealed system messages within LLMs as a preventive measure against jailbreak exploits. Through systematic exploration, we seek to shed light on this overlooked aspect of LLM security, paving the way for more effective mitigation strategies against jailbreaks.
To sum up, the main contributions of the paper are as follows:
@@
  • We introduce a novel perspective by examining the role of system messages in LLM security and jailbreak research.
  • We first conduct comprehensive tests on the impact of system messages, answering the question: "Is the system message really important for jailbreaks in LLMs?" By analyzing the influence of varying system messages on the transferability of jailbreak prompts, we identify a general relationship between the system message and jailbreak rates. In addition, we answer the question of whether it is possible to improve the robustness of LLMs to jailbreaks by making minor changes in the system message.
@@
  • To enhance resilience against jailbreak prompts, we investigate the optimization of system messages. We employ evolutionary algorithms to optimize the system messages, as most commercial LLMs can only be accessed through APIs. Specifically, in the case of minor changes in system messages, we propose System Message Evolutionary Algorithm (SMEA): SMEA-R, SMEA-C, and SMEA-X, each tailored to generate diverse and highly secure system messages. Through experiments, we validate the efficacy of SMEA, demonstrating its robustness to jailbreak prompts. This approach not only strengthens LLM security but also raises the bar for jailbreak prompts, contributing to the advancement of research in LLM security.
@@ 2 Background Knowledge
In this section, we aim to provide a comprehensive overview of the background knowledge. We begin by introducing related information about LLMs. Subsequently, we provide a detailed exposition on jailbreak prompts, shedding light on their significance. Lastly, given the incorporation of evolutionary algorithms in our study, we provide a general overview of how evolutionary algorithms leverage populations to iteratively obtain high-quality solutions.
@@ 2.1 Large Language Models
Large Language Models (LLMs), characterized by their extensive parameter count and training on vast corpora of human natural language, have emerged as a cornerstone in deep learning frameworks Chang et al. (2023). With parameter counts reaching into the billions, LLMs possess language capabilities akin to human fluency.
Presently, mainstream LLMs such as ChatGPT OpenAI (2022), GPT4 Bubeck et al. (2023), LLAMA2 Touvron et al. (2023) are founded on the transformer decoder architecture, operating in an autoregressive manner. This operational mode enables these models to predict subsequent words in a sequence based on preceding input text. Upon receiving input $\{w_{1},w_{2},\cdots,w_{t}\}$, LLMs predict the most probable next word $w_{t+1}$, subsequently incorporating it into the input sequence $\{w_{1},w_{2},\cdots,w_{t},w_{t+1}\}$ for further prediction. This iterative process continues until an end-of-sequence token is reached or the maximum sequence length supported by the LLMs is attained.
In developer mode, when interfacing with LLMs via the API, accessing to system messages is facilitated. These system messages typically exist in both long, short and no formats, as outlined in Appendix A.3. Conventionally, system messages serve as a protective barrier to reinforce security of LLMs. As shown in Figure LABEL:jail_example (b) (c), the LLM cannot output harmful content when just changing the system message. Despite their presumed significance, systematic investigations into system messages remain scarce. Hence, our research endeavors to bridge this gap by conducting a comprehensive exploration of system messages.
@@ 2.2 Jailbreak Prompts
Jailbreak prompts are meticulously constructed inputs engineered to provoke LLMs into generating harmful content Wei et al. (2023); Liu et al. (2023). Illustrated in Figure LABEL:jail_example (b), we provide an example of a jailbreak prompt capable of coaxing GPT3.5-turbo-0613 OpenAI (2022) with a short system message to produce deleterious output, with the red segments representing malicious queries. Despite concerted efforts to bolster LLMs’ resistance against such prompts Bai et al. (2022a); Ouyang et al. (2022); Christiano et al. (2017); Bai et al. (2022b), existing research suggests persistent vulnerabilities Zou et al. (2023); Huang et al. (2023); Lapid et al. (2023); Carlini et al. (2023); Qi et al. (2023); Chao et al. (2023); Yu et al. (2023); Li et al. (2023). For open source LLMs, Zou et al. (2023) proposed a novel jailbreak approach named GCG by appending suffixes to malicious questions, leveraging the autoprompt Shin et al. (2020) concept for open-source models, while also elucidating the transferability of jailbreak prompts. Huang et al. (2023) underscored the profound impact of parameter configurations on LLMs vulnerabilities. For black box LLMs, Lapid et al. (2023) employed genetic algorithms to craft adversarial suffixes for breaching black-box LLMs. However, the above mentioned methods of adding suffixes produces unnatural jailbreak prompts and more research focused on generating natural jailbreak prompts. Yu et al. (2023) advocated for applying software testing principles to LLMs jailbreaking, while Chao et al. (2023) proposed using one LLM to jailbreak another. Li et al. (2023) engineered a multi-step attack against ChatGPT, targeting the extraction of sensitive personal data, thereby raising profound privacy concerns and significant security risks for LLMs applications. In addition to the above recent investigations, Carlini et al. (2023) have demonstrated that simply manipulating image inputs can compromise the security of multi-modal models. Furthermore, research Qi et al. (2023) suggests that fine-tuning aligned LLMs on adversarial examples can jeopardize secure alignment. There is no doubt that jailbreak prompts pose a significant obstacle to the widespread use of LLMs.
@@ 2.3 Evolutionary Algorithms
Evolutionary algorithms Eiben et al. (2015) renowned as heuristic search techniques, using the principle of natural selection to optimize populations. Prominent algorithms in this domain include Differential Evolution (DE) Storn and Price (1997), NSGA-II Deb et al. (2002), IBEA Zitzler and Künzli (2004), and MOEA/D Zhang and Li (2007), all of which encompass three fundamental operators: crossover, mutation, and selection.
@@
    Crossover operator: This operator selects two individuals from the population and combines their information to produce offspring. This process is repeated until the desired number of offspring is generated, maintaining the population size.
@@
    Selection operator: Following mutation, the new population is merged with the previous population. Individuals are selected based on their fitness, with the best-performing individuals being chosen to form the new population for the next iteration.
Through successive iterations of these operations, individuals with inferior performance are gradually replaced, leading to the emergence of optimal and diverse individuals in the final population.
@@ 3.1 Models and Datasets
We select GPT3.5-turbo-0613 OpenAI (2022), LLAMA2 (7b, 7b-chat, 13b, 13b-chat) Touvron et al. (2023), VICUNA [^1] In this paper, VICUNA referred to exclusively denotes the version VICUNA-v1.5. (7b, 13b) Chiang et al. (2023), as LLMs to investigate the impact of system messages on jailbreaks. It’s worth noting that the GPT3.5-turbo-0613, LLAMA2-7b-chat, LLAMA2-13b-chat have undergone RLHF Bai et al. (2022a); Ouyang et al. (2022); Christiano et al. (2017); Bai et al. (2022b) techniques. To systematically explore this impact, we utilize GPTFuzzer Yu et al. (2023) to generate jailbreak prompts from a collection of $77$ prompt templates Liu et al. (2023) and $100$ questions Bai et al. (2022a); Liu et al. (2023). Specifically, we employ GPTFuzzer to generate $300$ jailbreak prompts on GPT3.5-turbo-0613, LLAMA2 (7b, 7b-chat, 13b, 13b-chat), VICUNA (7b, 13b) under three conditions respectively: with the short system message, the long system message, and no system message. Details regarding these system messages are provided in Appendix A.3. For clarity, we refer to jailbreak Prompts dataset from models with Short system messages as PS, jailbreak Prompts dataset from models with Long system messages as PL, and jailbreak Prompts dataset from models with the No system messages as PN.
@@ 3.2 Evaluation
In our study, we evaluate the success of LLMs jailbreaks using the Attack Success Rate (ASR) metric Zou et al. (2023); Lapid et al. (2023); Qi et al. (2023); Google (2023b); Chao et al. (2023); Huggingface (2023a); Rae et al. (2021). Previous studies have employed various criteria to evaluate the success of jailbreaks, which we summarize as follows:
  • Sub-string Matching: Researchers determined whether a jailbreak has occurred by identifying the presence of certain string in the output of LLMs Zou et al. (2023); Lapid et al. (2023). This evaluation is efficient, although it suffered from coarse-grained assessments and susceptibility to errors.
  • Toxicity: In addition, some researchers Qi et al. (2023) employed a rater such as perspective API Google (2023b) to evaluate the toxicity of responses, categorizing prompts as jailbreak prompts according to the toxic score of responses.
  • LLMs Evaluation: Moreover, some researchers used LLMs to evaluate responses Chao et al. (2023); Markov et al. (2023), with strict criteria. Some explored leveraging the few-shot learning capabilities of LLMs such as ChatGLM Huggingface (2023a) for evaluation Wang et al. (2020). However, existing studies Zhao et al. (2021); Lu et al. (2021); Ma et al. (2023) have emphasized the substantial impact of example selection and ordering on performance, which could potentially lead to outcomes close to randomness. Another way is fine-tuning Gopher Rae et al. (2021) in Bot-Adversarial Dialogue (BAD) Xu et al. (2021) for evaluation. However, using LLMs for evaluation runs too slow. However, the cost of using LLMs for evaluation is too high.
  • RobertaYu et al. (2023) evaluated responses using Roberta Liu et al. (2019) trained on the manually datasets human judgments and they compared their method with previous method. Their evaluation metrics includes True Positive Rate (TPR), False Positive Rate (FPR), and time efficiency. Additionally, they found that the Roberta model demonstrates promising performance in these metrics.
To sum up, we adopt the evaluation methodology proposed by Yu et al. (2023) to determine whether LLMs have exhibited jailbreak behaviors.
4 Jailbreak experiments with different system messages
To determine whether the system message is crucial for jailbreaks in LLMs, in this section we perform the following experiments.